Circular No. 8755

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 S5 (HILL) R. E. Hill reports his discovery of a comet on images taken in the course

of the Catalina Sky Survey with the 0.68-m Schmidt telescope (discovery observation tabulated below), the object described as having a tail ∼ 10′′ long to the southwest.

2006 UT  α_{2000}  δ_{2000}  Mag.
Sept.28.28000  0^{h}50^{m}22.^{s}81  +19^{o}01′40′′.8  18.6

Additional astrometry reported following posting on the ‘NEO Confirmation Page’, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2006-S84.

T = 2007 Dec. 24.713 TT  ω = 183.^{o}411
Ω = 283.471 }2000.0
q = 2.54479 AU  i = 10.189

COMET P/2006 S6 (HILL) Hill reports his discovery of another comet on images taken with the

Catalina 0.68-m Schmidt telescope, this one having a tail ∼ 10′′–15′′ long to the west-southwest.

2006 UT  α_{2000}  δ_{2000}  Mag.
Sept.28.40022  3^{h}04^{m}27.^{s}33  +25^{o}59′37′′.4  18.7

Additional astrometry reported following posting on the ‘NEO Confirmation Page’ [including prediscovery observations from Catalina (Aug. 29, Sept. 18) and from LINEAR (Sept. 18)], the following elliptical orbital elements, and an ephemeris appear on MPEC 2006-S85.

T = 2006 Oct. 18.054 TT  ω = 31.^{o}181
e = 0.42491  Ω = 9.070 }2000.0
q = 2.39862 AU  i = 13.178
a = 4.17087 AU n^{o} = 0.115708 P = 8.52 years

SUPERNOVAE 2006gi IN NGC 3147 G. Duszanowicz, Akersberga, Sweden, provides the following predis-

covery unfiltered CCD red magnitudes of 2006gi (cf. IAUC 8751): Aug. 31.89 UT, 18.1; Sept. 12.88, 16.1; 14.97, 16.0.

2006 September 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green